# Learning UAV-based path planning for efficient localization of objects using prior knowledge

### v1

@@ 3.3 Experiment 3: Influence of prior knowledge quality
Figure 13 shows the relation between flight path length and the number of found objects for different levels of prior knowledge quality. When having a perfect prior knowledge map without mistakes, the DQN learned a policy that found all objects. Even for prior knowledge of low quality, the DQN learned a policy to quickly find most objects. However, not all objects were found before the battery was empty. The higher the quality of the prior knowledge, the more objects were found. In complete absence of prior knowledge (level none) the agent still found some objects, however, a row-by-row flight path is more efficient in that case. The levels moderate and above outperformed the baseline by having found more than 80% of the objects in 71, 69 and 67 steps respectively compared to 206 steps for the row-by-row flight path. Level low needed around 300 steps to find 80% of the object.
Figure 13: Effect of the different prior knowledge quality levels on the percentage of found objects for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Figure 14 shows some flight paths for the DQN agents with different levels of prior knowledge quality. Without prior knowledge (Figure 14a), the agent quickly found the objects around the start location of the drone based on information from the local map, but failed to find objects further away from the start location. When having perfect prior knowledge (Figure 14e), the agent quickly finds all objects without wandering around.
Figure 14: Single flight paths of the RL agent for prior knowledge quality level none (no prior knowledge) (a), low (b), moderate (c), high (d) and perfect (e) and the baseline row-by-row flight path (f). The detected objects are indicated with black dots and the not detected objects with gray dots. The red line indicates the flight path till 80% of the objects are found, the blue line the complete flight path till all objects were found or the battery was empty.
@@ 3.4 Experiment 4: Effect of different stopping criteria
Figure 15 shows the percentage of objects found and flight path length using different stopping criteria. Setting a threshold on coverage resulted in a high percentage of found objects, but also a long path length. This indicates that the agent had difficulty fulfilling the coverage threshold before the battery was empty. Stopping the search task when there were no new detections during 15, 25 or 50 consecutive steps did result in a significantly shorter path length. However, terminating the task 15 or 25 steps after the last detection was too soon, indicated by the lower percentage of found objects. Terminating the search task when there were no new objects detected in the previous 50 steps resulted in a high number of found objects and a shorter flight path than the default stopping criterion (stop searching when all objects were found). Using a learned land action that terminates the search yielded a high percentage of found objects and a very short path length. When it is not essential to find all objects and a short flight path length is important, using a learned land action to terminate the search is suitable.
Figure 15: Boxplot showing the effect of different stopping criteria on the percentage of found objects and path length (right axis).
@@
Figure 16 shows the distribution of the action value of the learned land action (after the softmax layer) compared to the percentage of found objects. As was expected, the number of times the land action got a high action value increased when the percentage of found objects increased, indicating that the agent learned that landing is only profitable after finding most objects. In 2% of the episodes, the drone landed directly without taking any flight actions.
@@
Figure 16: Histogram showing the relation between the action value after the softmax layer for the landing action and the percentage of found objects. The colors are normalized column-wise.
@@ 4 Discussion
The results showed that, with prior knowledge, the RL agent was able to find a shorter path than a baseline row-by-row flight path for finding the objects when they were non-uniformly distributed. Even with a low-quality prior knowledge and high number of detection errors, the learned search policy still outperformed a baseline row-by-row flight path in terms of path length and was able to find most objects. Especially when the objects are not uniformly distributed, there is great potential in a learned search policy over a traditional row-by-row flight path if a perfect mapping of objects is not required. When there was no perfect prior knowledge available, the RL agent still found most objects quicker than the row-by-row flight path, but was not able to find all objects. In section 4.1 we discuss the action selection by the agent, in section 4.2 the used action space, section 4.3 learning in absence of prior knowledge, section 4.4 the application and finally, section 4.5 discusses the influence of the assumptions on the applicability in the real world.
@@ 4.1 Action selection
Towards the end of some episodes, the agent sometimes got stuck in the same spot when not all objects were found by repeating the same two actions, such as ’fly north’ and ’fly south’ or ’fly east’ and ’fly west’ (Figure 14b for an example). Using a learned land action in experiment 4 solved this issue as it allowed the RL agent to decide when to terminate the search task and thus, when it was not profitable anymore to search for the last objects. However, we observed some cases where the drone directly terminated the search task. This only happened in 2% of the episodes, in contrast to Yang et al. (2018) and Druon et al. (2020). Yang et al. (2018) indicated a 20-30% lower performance when using a stopping action and Druon et al. (2020) attributed around 50 percent of the failure cases to selecting the stopping action at the wrong moment. The difference in performance may be due to their use of an end-to-end network architecture that takes images as input, unlike the abstract representation used in this work. As these images have a larger variability than the abstract representation we use, the network is likely to be more uncertain about its predictions, which makes it more profitable to terminate earlier. To reduce the remaining 2% of direct terminations, a penalty for missing objects when landing could be introduced during training. This would encourage the agent to explore at least parts of the field before landing.
@@ 4.2 Action space
Similar to most research on RL-based path planning for UAVs (e.g. Gugan and Haque (2023); Husnain et al. (2023)), in this work the environment is simplified into a 2D representation, where the UAV flies at a fixed altitude. To further increase the efficiency of the RL learned policy, the environment could be changed to 3D by involving actions for altitude changes, which can have several advantages for a search policy. In context of object search, it can be used to get a (low resolution) overview of the whole field at a high altitude, which can be used as prior knowledge in the global map. The drone can then learn to fly lower to inspect parts of the field in high resolution.
4.3 Learning in absence of prior knowledge
Experiment 3 showed that learning a search policy without any prior knowledge remained a challenge. Because the agent had no clue where the objects were, gathering useful state transition vectors for the experience replay buffer was difficult, which prevented the agent from learning useful actions. Possibly, curriculum learning can be used to gradually build up the challenge for the agent (Narvekar et al., 2020), starting with a simulation that has full prior knowledge and slowly building up to a simulation without prior knowledge.
@@ 4.4 Application
The usability of the presented learned policy depends on the type of application. When the goal of an application is to find most objects in a short period of time, the learned policy clearly outperforms the baseline row-by-row flight path. However, there is no guarantee that all objects are found, which limits the usability of such learned policy in applications where it is crucial to find all objects. For the use-case of weed detection in arable fields, it may be fine to detect most weeds and miss some. Since weed detection requires multiple flights throughout the season, weeds that were not detected this flight can be detected in a later flight.
@@ 4.5 Impact of assumptions on applicability in the real world
Although the presented work was based on a abstracted simulated environment, the proposed method can easily be applied to a real-world scenario with a few additions. The simulator was designed to be used in combination with an object-detection system, prior knowledge information and a flight controller (Figure 1). With a trained detector, objects can be detected in a real camera image, which results in the object map input to the RL. The resulting discrete flight actions can executed on a real drone using the drone’s flight controller. Below we discuss the underlying assumptions of the simulation and RL policy, as mentioned in the introduction and discuss the implications for real-world use:
@@
Drone can accurately execute flight actions: the flight controller of the drone should be able to translate the discrete flight actions into motor signals. The discrete actions are in four directions with a distance determined by the grid-cell size. The localization accuracy of the drone should therefore be higher than the grid-cell size, to be able to accurately move the drone to a specific grid-cell. When using grid-cells of, for example, 1x1m, a drone equipped with RTK-GPS (accuracy of around 2-3cm) is more than accurate enough to execute the RL policy. As the use of RTK-GPS is quite standard on agricultural drones, this assumption can be met.
@@
Field size: since the prior knowledge map is related to the field size and the Q-network uses a combination of convolutions and fully-connected layers (Figure 4), a larger field increases the size of the flattened layer and thereby the number of trainable parameters. This will require more training iterations. Experiment 3 showed that the number of found objects only dropped at a very low prior knowledge quality (and thus very low resolutions), which indicates that the resolution of the prior knowledge may be low. Therefore, larger fields can be used with the same number of trainable parameters by applying a larger down-sampling factor (increasing $g_{\textrm{global}}$), or by incorporating prior knowledge in a different way.
@@
Prior knowledge of field boundaries: field boundaries are static and available on forehand. In this work, we limited the experiments to squared fields. Although not specifically tested, applying the learned search policy in non-squared fields can be achieved by changing the field-area map containing information if the cell is inside or outside the field. Additionally, obstacles within the field, such as trees, can be avoided by mapping them as non-field area. As the violation of the field-area map results in a negative reward, it is very likely that the RL agent will learn to deal with it. Work of Theile et al. (2021) showed that a RL-learned policy with a comparable architecture is able to avoid buildings in a city when visiting specific locations using drones by encoding these positions in a no-fly-zone, which is in principle comparable with our field-area map.
@@
A decently trained detection network: the results of experiment 2 showed that the RL agent is robust to errors in the object detection. Hence, the detection network does not have to be very accurate. Even when the detection network generates a lot of false positives, the RL policy was able to find most of the true positives. Typical object detection performance in agriculture lay around 0.7-0.9 F1-score (Rai et al., 2023; Ruigrok et al., 2023; Rehman et al., 2024). This roughly corresponds with level ’high’ in experiment 2, which indicates that the RL agent can deal with input from a real field.
@@
Prior knowledge of location of objects: although some level of prior knowledge is required, it does not need to be very detailed. Experiment 3 showed that the prior knowledge did not need to be very detailed or very accurate. This prior knowledge can be created by, for example, using the detections from a few higher altitude images, data from previous flights, or other data sources such as satellite images. For many agricultural tasks, like weed detection and disease detection, such rough prior knowledge is available. For instance, many weeds appear in the same area in subsequent years and spread very locally. Information from previous year can then be used as prior knowledge.
@@
The above analysis indicates that all the underlying assumptions of the abstracted RL agent and simulation environment are not problematic for a real-world application of a RL learned search policy. Especially in use-cases where the target objects are distributed non-uniformly in the field and when it is not crucial to find all objects, there is much to gain when specifically searching for objects instead of flying over and covering the whole field.

### v2

@@ 3.3 Experiment 3: Influence of prior knowledge quality
Figure 14 shows the relation between flight-path length and the number of found weeds for different levels of prior knowledge quality. When having a perfect prior knowledge map without mistakes, the DQN learned a policy that found all weeds. Even with prior knowledge of low quality, the DQN learned a policy to quickly find most weeds. However, not all weeds were found before the battery was empty. The higher the quality of the prior knowledge, the more weeds were found. In complete absence of prior knowledge (level none) the agent still found some weeds, however, a row-by-row flight path is more efficient in that case. The levels moderate and above outperformed the baseline by finding more than 80% of the weeds in 71, 69, and 67 steps, respectively, compared to 206 steps for the row-by-row flight path. Level low needed around 300 steps to find 80% of the weed.
Figure 14: Effect of the different prior knowledge quality levels on the percentage of found weeds for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Table 7: Percentage of found weeds at 100, 200, and 300 flight steps for different prior knowledge quality levels for the policy learned by DQN, and the baseline row-by-row flight path. The values show the mean ± the standard deviation. Values indicated with a ’*’ have a significant ($\alpha=0.001)$ higher percentage of found weeds than the baseline row-by-row flight path (Welch’s t-test).
Method 100 steps 200 steps 300 steps
row-by-row 0.37±0.30 0.76±0.25 1.00±0.00
dqn-none 0.35±0.32 0.45±0.33 0.49±0.33
dqn-low 0.72±0.24* 0.78±0.22 0.80±0.20
dqn-moderate 0.90±0.12* 0.96±0.05* 0.97±0.05
dqn-high 0.92±0.10* 0.97±0.03* 0.98±0.03
dqn-perfect 0.94±0.08* 0.99±0.02* 0.99±0.01
Table 7 shows the percentage of found weeds for the different levels of prior knowledge corresponding to 100, 200, and 300 flight steps for the DQN policy and the baseline row-by-row flight path. Levels ’moderate’ and higher had a significantly higher number of found weeds than the baseline at 200 flight steps. Level ’low’ only outperformed the baseline at 100 flight steps. Without prior knowledge, the row-by-row flight path was more efficient even at a low number of flight steps. At 300 flight steps, the baseline found all weeds, compared to 97%, 98% and 99% for the levels ’moderate’, ’high’, and ’very-high’ respectively.
Figure 15 shows some flight paths for the DQN agents with different levels of prior knowledge quality. Without prior knowledge (Figure 15a), the agent quickly found the weeds around the start location of the drone based on information from the local map, but failed to find weeds further away from the start location. When having perfect prior knowledge (Figure 15e), the agent quickly finds all weeds without wandering around.
Figure 15: Single flight paths of the RL agent for prior knowledge quality level none (no prior knowledge) (a), low (b), moderate (c), high (d), and perfect (e), and the baseline row-by-row flight path (f). The detected weeds are indicated with black dots and the undetected weeds with gray dots. The red line indicates the flight path till 80% of the weeds are found, the blue line the complete flight path till all weeds were found or the battery was empty.
@@ 3.4 Experiment 4: Effect of different stopping criteria
Table 8 shows the percentage of weeds found and flight-path length using different stopping criteria. Setting a threshold on coverage resulted in a high percentage of found weeds, but also a long path length. This indicates that the agent had difficulty fulfilling the coverage threshold before the battery was empty. Stopping the search task when there were no new detections during 15, 25, or 50 consecutive steps resulted in a significantly shorter path length than the baseline row-by-row flight path. However, terminating the task 15 or 25 steps after the last detection was too soon, indicated by the lower percentage of found weeds. Terminating the search task when there were no new weeds detected in the previous 50 steps resulted in a high number of found weeds and a shorter flight path than the default stopping criterion (stop searching when all weeds were found). Using a learned land action that terminates the search yielded a high percentage of found weeds and a very short path length. Compared to the baseline row-by-row flight path, it yielded a 74% shorter flight path at the cost of a 12% lower percentage of found weeds. When it is not essential to find all weeds and a short flight-path length is important, using a learned land action to terminate the search is suitable.
Table 8: Percentage of found weeds and path length for the different stopping criteria and the baseline row-by-row flight path. The values show the mean ± the standard deviation. Values indicated with a ’*’ have a significantly shorter ($\alpha=0.001)$ path length than the baseline row-by-row flight path (Welch’s t-test).
Method Percentage found weeds Path length
row-by-row 1.00±0.00 276±0
all weeds (default) 0.98±0.03 281±130
coverage 50% 0.94±0.08 239±97*
coverage 75% 0.97±0.06 371±14
15 steps no new weeds 0.45±0.42 52±40*
25 steps no new weeds 0.72±0.38 98±52*
50 steps no new weeds 0.94±0.14 193±79*
land action 0.88±0.16 71±31*
@@
Figure 16 shows the distribution of the action value of the learned land action (after the softmax layer) compared to the percentage of found weeds. As was expected, the number of times the land action got a high action value increased when the percentage of found weeds increased, indicating that the agent learned that landing is only profitable after finding most weeds. In 2% of the episodes, the drone landed directly without taking any flight actions.
@@
Figure 16: Histogram showing the relation between the action value after the softmax layer for the landing action and the percentage of found weeds. The colors are normalized column-wise.
3.5 Experiment 5: Transferability of simulation-learned policy to a real-world application
Figure 17 shows the percentage of found weeds against the flight-path length on the four real-world datasets for the DQN policy and the baseline row-by-row flight path. Till 200 flight steps, the DQN policy found more weeds than the baseline. When the drone landed, the percentage of found weeds on the real-world image data ranged between 68% to 97% for the DQN policy. Due to detection errors, the baseline did not find all weeds, with the total percentage of found weeds ranging between 82% to 97%. On average, the DQN policy found $81.2\pm 10.4\%$ of the weeds before landing in $94\pm 31$ flight steps, whereas the row-by-row flight path found $90.9\pm 5.8\%$ of the weeds in $276\pm 0$ flight steps. This corresponds to a 66% shorter flight path compared to the row-by-row flight path at the cost of a 10% lower percentage of found weeds.
Figure 17: Percentage of found weeds and the flight-path length for the policy learned by DQN and the baseline row-by-row flight path on the four real-world datasets.
Figure 18 shows the flight path of the DQN policy and the baseline row-by-row flight path for real-world datasets. On dataset 1, the DQN policy missed a cluster of weeds, which explains the lower percentage of found weeds for dataset 1 in Figure 17. On datasets 2 and 4, the DQN policy found all clusters, but missed some weeds within the clusters. Both the DQN policy and the baseline made some false-positive and false-negative detections.
Figure 18: Flight path on dataset 1 (a,e), 2 (b,f), 3 (c,g) and 4 (d,h) for the DQN policy (a–d) and the baseline row-by-row flight path (e–h). Blue, orange, and red markers indicate a true positive, false positive, and false negative object detections, respectively. The start and end position of the UAV is indicated by an arrow.
@@ 4 Discussion
The results showed that, with prior knowledge, the RL agent was able to find a shorter path than a baseline row-by-row flight path for finding the weeds when they were non-uniformly distributed. Even with low-quality prior knowledge and a high number of detection errors, the learned search policy still outperformed a baseline row-by-row flight path in terms of path length and was able to find most weeds. Especially when the weeds are not uniformly distributed, there is great potential in a learned search policy over a traditional row-by-row flight path if a perfect mapping of weeds is not required. When there was no perfect prior knowledge available, the RL agent still found most weeds quicker than the row-by-row flight path, but was not able to find all weeds. The simulation-learned policy is transferable to a real-world application, as was shown in experiment 5. Section 4.1 discusses the DQN policy results compared to the literature, section 4.2 the action selection by the agent, section 4.3 the used action space, section 4.4 learning in absence of prior knowledge, section 4.5 the application and finally, section 4.6 discusses the influence of the assumptions on the applicability in the real world.
4.1 Comparison with literature
All results in this paper were compared to a baseline row-by-row flight path, which is most commonly used in practice. On real-world data, the DQN policy yielded a 66% shorter flight path at the cost of a 10% lower percentage of found weeds. Several alternative UAV path planning methods for object detection in fields are proposed in literature. For instance, van Essen et al. (2025) showed a rule-based adaptive path planner that resulted in a 37% shorter flight path at the cost of a 2% lower F1-score compared to a row-by-row flight path. Popovic et al. (2017) introduced a path planner for active weed classification that found 85% of the weeds in 150 seconds, whereas a row-by-row flight path took 300 seconds to find the same number of weeds. Both papers use different data, which makes a direct comparison impossible, but they do show the same trade-off between accuracy and flight path length. Similarly to the results in this paper, they both show that the flight paths were more efficient when the objects were non-uniformly distributed.
@@ 4.2 Action selection
Towards the end of some episodes, the agent sometimes got stuck in the same spot when not all weeds were found by repeating the same two actions, such as ’fly north’ and ’fly south’ or ’fly east’ and ’fly west’ (Figure 15b for an example). Using a learned land action in experiment 4 solved this issue as it allowed the RL agent to decide when to terminate the search task, and thus, when it was not profitable anymore to search for the last weeds. However, we observed some cases where the drone directly terminated the search task. This only happened in 2% of the episodes, in contrast to Yang et al. (2018) and Druon et al. (2020). Yang et al. (2018) indicated a 20-30% lower performance when using a stopping action, and Druon et al. (2020) attributed around 50 percent of the failure cases to selecting the stopping action at the wrong moment. The difference in performance may be due to their use of an end-to-end network architecture that takes images as input, unlike the abstract representation used in this work. As these images have a larger variability than the abstract representation we use, the network is likely to be more uncertain about its predictions, which makes it more profitable to terminate earlier. To reduce the remaining 2% of direct terminations, a penalty for missing weeds when landing could be introduced during training. This would encourage the agent to explore at least parts of the field before landing.
@@ 4.3 Action space
Similar to most research on RL-based path planning for UAVs (e.g. Gugan and Haque (2023); Husnain et al. (2023)), in this work, the environment is simplified into a 2D representation, where the UAV flies at a fixed altitude. To further increase the efficiency of the RL learned policy, the environment could be changed to 3D by involving actions for altitude changes, which can have several advantages for a search policy. In the context of weed detection, it can be used to get a (low resolution) overview of the whole field at a high altitude, which can be used as prior knowledge in the global map. The drone can then learn to fly lower to inspect parts of the field in high resolution.
4.4 Learning in the absence of prior knowledge
Experiment 3 showed that learning a search policy without any prior knowledge remained a challenge. Because the agent had no clue where the weeds were, gathering useful state transition vectors for the experience replay buffer was difficult, which prevented the agent from learning useful actions. Possibly, curriculum learning can be used to gradually build up the challenge for the agent (Narvekar et al., 2020), starting with a simulation that has full prior knowledge and slowly building up to a simulation without prior knowledge.
@@ 4.5 Application
The use case in this paper is weed detection in agricultural fields. However, the learned planner could be applied to any task where the target objects are non-uniformly distributed in the environment and the goal of the application is to quickly find most of them. For these applications, the learned policy clearly outperforms the baseline row-by-row flight path. However, there is no guarantee that all objects are found, which limits the usability of such a learned policy in applications where it is crucial to find all objects. For the use case of weed detection in arable fields, it may be fine to detect most weeds and miss some. Since weed detection requires multiple flights throughout the season, weeds that were not detected in this flight can be detected in a later flight.
The agent was trained on a large variety of different simulated object distributions. The trained agent may work for real-world applications where the number of clusters, the number of objects, and the spatial distribution within the clusters fall within the randomized ranges used during training. However, when there are large differences between the real-world distribution of the weeds and the distribution in the training simulation, changing the simulation parameters and retraining the agent may be required.
During a real-world application, the inference time equals the sum of both the detection network and the DQN. On embedded devices, such as a NVIDIA Jetson Orin, the used detection network, YOLOv8-nano, has an inference time of around 20 ms (Rey et al., 2025). Since the DQN is even smaller, running both the object detection network and the DQN on board a drone in real time should be feasible. Future work will focus on applying the simulation-trained RL policy to a real drone on a real-world application.
@@ 4.6 Impact of assumptions on applicability in the real world
Although most of the presented work was based on an abstracted simulated environment and pre-recorded datasets, the proposed method can easily be applied to a real-world scenario with a few additions. The simulator was designed to be used in combination with an object-detection system, prior knowledge information, and a flight controller (Figure 1). Using a trained object detector, weeds can be detected from a camera image to generate the local weed map as input to the RL, as demonstrated in experiment 5. The resulting discrete flight actions can be executed on a real drone using the drone’s flight controller. Below, we discuss the underlying assumptions of the simulation and RL policy, as mentioned in the introduction, and discuss the implications for real-world use:
@@
Drone can accurately execute flight actions: the flight controller of the drone should be able to translate the discrete flight actions into motor signals. The discrete actions are in four directions with a distance determined by the grid-cell size. Due to inaccuracies in the drone’s position and attitude, the estimated location of the drone is imperfect. The localization accuracy of the drone should therefore be higher than the grid-cell size, to be able to accurately move the drone to a specific grid-cell. Although not tested in this paper, when using grid-cells of, for example, 1x1 m, a drone equipped with RTK-GPS (accuracy of around 2-3cm) should be more than accurate enough to execute the RL policy. As the use of RTK-GPS is quite standard on agricultural drones, this assumption can be met.
@@
Field size: since the prior knowledge map is related to the field size and the Q-network uses a combination of convolutions and fully-connected layers (Figure 5), a larger field increases the size of the flattened layer and thereby the number of trainable parameters. This will require more training iterations. Experiment 3 showed that the number of found weeds only dropped at a very low prior knowledge quality (and thus very low resolutions), which indicates that the resolution of the prior knowledge may be low. Therefore, larger fields can be used with the same number of trainable parameters by applying a larger down-sampling factor (increasing $g_{\textrm{global}}$), or by incorporating prior knowledge in a different way.
@@
Prior knowledge of field boundaries: field boundaries are static and available on forehand. In this work, we limited the experiments to squared fields. Although not specifically tested, applying the learned search policy in non-squared fields can be achieved by changing the field-area map, containing information on whether the cell is inside or outside the field. Additionally, obstacles within the field, such as trees, can be avoided by mapping them as non-field areas. As the violation of the field-area map results in a negative reward, it is very likely that the RL agent will learn to deal with it. Work of Theile et al. (2021) showed that an RL-learned policy with a comparable architecture is able to avoid buildings in a city when visiting specific locations using drones by encoding these positions in a no-fly-zone, which is in principle comparable with our field-area map.
@@
A decently trained detection network: due to factors such as variation in the natural environment and motion blur, the output of an object detection network contains errors. The results of experiment 2 showed that the RL agent is robust to these errors in the object detection. Hence, the detection network does not have to be very accurate. Even when the detection network generates a lot of false positives, the RL policy was able to find most of the true positives. Typical object detection performance in agriculture lies around 0.7-0.9 F1-score (Rai et al., 2023; Ruigrok et al., 2023; Rehman et al., 2024). This roughly corresponds with level ’high’ in experiment 2, which indicates that the RL agent can deal with input from a real field. Experiment 5 confirmed this by finding most of the weeds in the real-world datasets using the RL agent in combination with a real object detection network.
@@
Some prior knowledge of location of objects: although some level of prior knowledge is required, it does not need to be very detailed. Experiment 3 showed that the prior knowledge did not need to be very detailed or very accurate, however, some prior knowledge was still required. In practice, this prior knowledge can be created by, for example, using the detections from a few higher altitude images (as done in experiment 5), data from previous flights, or other data sources such as satellite images. For many agricultural tasks, like weed detection and disease detection, such rough prior knowledge is available. For instance, the location of many weeds may be predicted by previous years’ location (Colbach et al., 2000), and spread very locally. Information from the previous year can then be used as prior knowledge.
@@
The above analysis indicates that all the underlying assumptions of the abstracted RL agent and simulation environment are not problematic for a real-world application of an RL learned search policy. Future research should focus on applying the DQN policy on a real drone to check the influence of the drone’s positional and attitude errors on the execution of the task. Especially in use-cases where the target objects are distributed non-uniformly in the field and when it is not crucial to find all objects, there is much to gain when specifically searching for objects instead of flying over and covering the whole field.